OpenStudy (eric_d):

The expression ax^3-8x^2+bx+6 is divisible by x^2-2x-3. Find values of a & b. @ikram002p

OpenStudy (ikram002p):

so (ax^3-8x^2+bx+6)= (ax-M) (x^2-2x-3) a.t M is devides 6 so it might be 1,2,3,6,-1,-2,-3,-6

OpenStudy (ikram002p):

so \((ax-M) (x^2-2x-3)=ax^3-2ax^2-3ax-Mx^3+2Mx+3M\)

OpenStudy (ikram002p):

\(ax^3-8x^2+bx+6 =ax^3-(M+2a)x^2+(2M-3a)x +3M\)

OpenStudy (ikram002p):

3M=6 so what is M value ?

OpenStudy (eric_d):

2

OpenStudy (ikram002p):

yes , so ur getting the qn ?

OpenStudy (ikram002p):

2M-3a=b M+2a=8

OpenStudy (eric_d):

M actually refers to...?

OpenStudy (ikram002p):

its a variable i added it so if \( ax^3-8x^2+bx+6 \) is divisible by \( x^2-2x-3 \) then \( ax^3-8x^2+bx+6= (\text{other term} )(x^2-2x-3)\) ok

OpenStudy (ikram002p):

(so other ) term can be (ax-M) we dnt know M value yet

OpenStudy (ikram002p):

so the equation will be like this \( ax^3-8x^2+bx+6 =(ax-M)( x^2-2x-3)\)

OpenStudy (ikram002p):

expand \((ax-M)( x^2-2x-3)=?\)

OpenStudy (eric_d):

Understood So, this's the method to solve this type of question..

OpenStudy (ikram002p):

yes ^^

OpenStudy (eric_d):

Another question...

OpenStudy (ikram002p):

wait but did u find a,b for this ?

OpenStudy (ikram002p):

sure ask ^^

OpenStudy (eric_d):

when polynomial x^2+ax+b is divided by (x-1) & (x+2), its remainder is 4 n 5 respectively. Find the values of a n b

OpenStudy (ikram002p):

so \(x^2+ax+b =(x-M)+\large \frac {4}{x-1}\) so \(x^2+ax+b =(x-N)+ \large \frac {5}{x-2}\)

OpenStudy (ikram002p):

its like 9/2 u can write the answer like 4+ 1/2 since

OpenStudy (ikram002p):

use long Division

OpenStudy (eric_d):

Okay, I'll try to solve this question using this method..

OpenStudy (ikram002p):

:D do u have other method ?

OpenStudy (eric_d):

Nope

OpenStudy (ikram002p):

if yes i would be happy to learn it :P

OpenStudy (ikram002p):

ok solve this , and tell me when u get it ^_^

OpenStudy (ikram002p):

sry i made a typo brb

hartnn (hartnn):

alternative method for the first problem, x^2-2x-3 =0 gives x= 3,-1 so, plug in x=3 in ax^3-8x^2+bx+6 =0 get one equation in a and b plug in x=-1 in ax^3-8x^2+bx+6 =0 get another equation in a and b solve simultaneously.

OpenStudy (eric_d):

This method seems to be much simpler

hartnn (hartnn):

alternative method for "when polynomial x^2+ax+b is divided by (x-1) & (x+2), its remainder is 4 n 5 respectively. Find the values of a n b" plug in x=1, in x^2+ax+b =4 get one equation in a and b plug in x=-2, in x^2+ax+b =5 get another equation in a and b solve simultaneously

hartnn (hartnn):

i have used remainder theorem .

OpenStudy (ikram002p):

i made a typo , i ment to write equation like this \(\large \frac{ x^2+ax+b}{x-1} =(x-M)+\large \frac {4}{x-1}\) \(\large \frac{ x^2+ax+b}{x-2} =(x-MN)+\large \frac {5}{x-2}\)

OpenStudy (ikram002p):

\(\large \frac{ x^2+ax+b}{x-2} =(x- N)+\large \frac {5}{x-2}\) *

OpenStudy (eric_d):

okay

OpenStudy (ikram002p):

yeah so first one it will be \(x^2+ax+b =(x-M ) (x-1) + 4\) \(x^2+ax+b =(x-N ) (x+2) + 5\)

OpenStudy (ikram002p):

^_^

OpenStudy (eric_d):

for the 2nd question.. a=2/3, b= 7/3 Another question: The expression x^3-4x^2+x+6 & x^3-3x^2+2x+k have common factor.Find possible values of k.

OpenStudy (eric_d):

Can I use remainder theorem to solve this question @hartnn

OpenStudy (anonymous):

too much work for the first question

OpenStudy (anonymous):

u can do it in 3 steps probably

hartnn (hartnn):

you can use remainder theorem there, but it won't help you find k

OpenStudy (ikram002p):

is it mono factor ?

OpenStudy (eric_d):

maybe

hartnn (hartnn):

heard about rational root theorem ?

OpenStudy (ikram002p):

yeah , that will work

OpenStudy (eric_d):

No.

hartnn (hartnn):

then which topic does this question belong to ?

OpenStudy (eric_d):

Chapter: Functions Sub topic is factor theorem.. I don't mind learning this rational root....

OpenStudy (ikram002p):

x^3-4x^2+x+6 =(x-a)(....) x^3-3x^2+2x+k =(x-a)(...) let (x-a ) be the common factor , then a should be a factor of both 6,k

OpenStudy (ikram002p):

factorize 6:- then possible values for a is \(\pm1,\pm 2 , \pm 3\)

OpenStudy (ikram002p):

but k= n a mmmm we need to find a upper bound for k

OpenStudy (ikram002p):

would u plz factorize this equation ? x^3-4x^2+x+6

hartnn (hartnn):

what i can think of as an alternative way is since we take 'x-a' as a factor, we can plug in x=a in both expressions and equate them a^3-4a^2+a+6 = a^3-3a^2+2a+k this gives you quadratic in a and k

hartnn (hartnn):

and since roots must be real you can do b^2-4ac >0 to find range of k

hartnn (hartnn):

but ultimately possible values of k will be from factors of 6 only.

OpenStudy (ikram002p):

\( x^3-4x^2+x+6 =(x+1)(x-2)(x-3)\) so three cases :- \( x^3-3x^2+2x+k =(x+1)(...)\) \( x^3-3x^2+2x+k =(x-2)(...)\) \( x^3-3x^2+2x+k =(x-3)(...)\)

OpenStudy (eric_d):

what must I do nxt

OpenStudy (ikram002p):

next make usfull of long Division

OpenStudy (ikram002p):



OpenStudy (eric_d):

Okay..will try now

OpenStudy (ikram002p):